# Ultraviolet photon production rates of the first stars: Impact on the He II 1640 Ă emission line from primordial star clusters and the 21-cm signal from cosmic dawn

### v1

@@ Abstract
The first stars, the chemically pristine Population III, likely played an important role in heating the intergalactic medium during the epoch of cosmic dawn. The very high effective temperatures ( K) predicted for the most massive Population III stars could also give rise to tell-tale signatures in the emission-line spectra of early star clusters or small galaxies dominated by such stars. Important quantities in modelling their observational signatures include their photon production rates at ultraviolet energies at which photons are able to ionize hydrogen and helium, dissociate molecular hydrogen and cause Lyman- heating.
Here, we model the spectral energy distributions of Population III stars to explore how these key quantities are affected by the initial mass and rotation of Population III stars given a wide range of models for the evolution of these stars. Our results indicate that rotating Population III stars that evolve to effective temperatures K could potentially give rise to a very strong HeII 1640 Ă emission line in the spectra from primordial star clusters, without requiring stellar masses of indicated by previous models for non-rotating Population III stars. At the same time, the observable impact on 21-cm signatures from cosmic dawn and epoch of reionization from our set of rotating stars that evolve to K are modest, and produce potentially detectable features in the global 21-cm signal and 21-cm power spectrum for high Population III star formation efficiencies only.
@@
Stars: Population III â stars: rotation â cosmology: dark ages, reionization, first stars â galaxies: high-redshift1 Introduction
@@
The first generation of stars, the chemically pristine Population III (hereafter Pop III), is believed to have played an important role in heating the early intergalactic medium during the epoch of cosmic dawn, in producing the first elements beyond lithium, and in the production of early black holes (for a recent review, see Klessen & Glover, 2023). Such stars probably started forming around redshift , about 100 Myr after the Big Bang, but may have continued to form in pockets of primordial gas throughout the reionization epoch and well into the post-reionization era (e.g. Johnson, 2010; Mebane et al., 2018; Liu & Bromm, 2020). In terms of observational probes, 21-cm cosmology may be able to probe the nature of Pop III stars at â30 through their impact on the temperature state of the intergalactic medium (e.g. Pochinda et al., 2024; Gessey-Jones et al., 2025; Ventura et al., 2025). At somewhat lower redshifts (â15) one may hope to detect Pop III signatures in individual objects, either among samples of gravitationally lensed, high-redshift stars (Windhorst et al., 2018; Zackrisson et al., 2024), in gravitationally lensed Pop III star clusters and galaxies (Zackrisson et al., 2012; Vanzella et al., 2023; Fujimoto et al., 2025), or within galaxies where Pop III star formation is able to continue alongside metal-enriched star formation (e.g. Sarmento et al., 2018; Venditti et al., 2023; Maiolino et al., 2024; Wang et al., 2024). Other ways of empirically studying the properties of Pop III stars include the chemical signatures of Pop III supernovae in gas absorption systems (e.g. DâOdorico et al., 2023; Sodini et al., 2024; Vanni et al., 2024), in high-redshift quasars (Yoshii et al., 2022) and in second-generation stars in the local Universe (e.g. Fraser et al., 2017; Ishigaki et al., 2018; Jiang et al., 2024), as well as the direct detection of supernovae (e.g. Moriya et al., 2022; Venditti et al., 2024), gamma-ray bursts (e.g. Lazar & Bromm, 2022), and tidal disruption events (Kar Chowdhury et al., 2024) produced by Pop III stars.
@@
Due to their chemically pristine composition, Pop III stars are expected to become more massive (e.g. Klessen & Glover, 2023), and reach higher main-sequence temperatures (e.g. Schaerer, 2002) than their metal-enriched descendants. This gives rise to higher hydrogen- and helium-ionizing photon production rates compared to metal-enriched stars, and also differences in the LymanâWerner (LW; -dissociating) and the Lyman-band rates important for the WouthuysenâField coupling during cosmic dawn. However, the production of photons in these different energy regimes depends not only on initial mass but also on stellar rotation (e.g. Yoon et al., 2012; Murphy et al., 2021b; Sibony et al., 2022; Liu et al., 2025a, b; Hawcroft et al., 2025), and simulations have indicated that at least some Pop III may rotate at close to break-up speed (Hirano & Bromm, 2018).
@@
Here, we use the Muspelheim (Zackrisson et al., 2024) models for the spectral energy distributions (SEDs) of Pop III stars to explore the impact of different sets of stellar evolutionary tracks on the ultraviolet (UV) radiation output of Pop III stars. These models use stellar atmosphere SEDs instead of the blackbody SEDs adopted in many previous studies (e.g. Yoon et al., 2012; Murphy et al., 2021b; Klessen & Glover, 2023). In section 2, we present these models and in section 3 we report the resulting hydrogen- and helium-ionizing photon production rates, the LW and the Ly-band rates, along with a comparison to results derived from blackbody SEDs. The impact on studies of the UV SEDs of Population III star clusters and galaxies is explored in Section 4 and the impact on 21 cm-cosmology is shown in Section 5. Various future avenues for improvement of the models are discussed in section 6. Section 7 summarises our findings.
@@ 2 SED models of Population III stars
The SED models for Pop III stars (Muspelheim v.1.0) presented by Zackrisson et al. (2024) couple sets of stellar evolutionary tracks to a grid of stellar atmospheres to produce SEDs as a function of Zero Age Main Sequence (ZAMS) mass and age. Important simplifications used in the Muspelheim v.1.0 models include the use of stellar atmospheres for spherical stars without mass loss, and with fixed primordial abundances of hydrogen and helium. Hence, these models neglect the non-sphericity that come from high rotation rates, the effects of winds, and the effects of surface helium enhancement and surface self-pollution of metals during the course of evolution. These shortcomings are further discussed in Section 6. In this paper, we make use of Muspelheim SED models based on selected Pop III tracks from Yoon et al. (2012) at â500 Mâ, Windhorst et al. (2018) at â100 Mâ, Murphy et al. (2021a) at â120 Mâ, Volpato et al. (2023) at â1000 Mâ and Costa et al. (2025) at â200 Mâ. Out of these sets, the Yoon et al. (2012) and Murphy et al. (2021a) sets feature models both with and without rotation. Although Yoon et al. (2012) present stellar evolutionary tracks for several different initial velocities in the range 0.0â0.6 (where is the Keplerian velocity at the equatorial surface), here we limit the discussion to the and 0.4 models, for simplicity.
@@
In this paper, we explore the UV properties of Pop III stars in terms of their production of H-, He- and He+-ionizing photons, as well as their production of LW and Lyman-band photons. The instantaneous photon production rates , expressed in units of photons per second, are defined as:
where is the monochromatic luminosity (erg-1 s-1 Ă -1) at wavelength , is Planckâs constant and the speed of light. The integration limits for the different rates are: and Ă for ; and Ă for ; and Ă for ; Ă and Ă for ; Ă and Ă for . We also present , the Lyman-band photon production count per baryon, integrated over the stellar lifetime , given by:
@@
3 Ionizing and LymanâWerner fluxes
@@
As previously shown by e.g. Bromm et al. (2001), Larkin et al. (2023), and Liu et al. (2025a), the SEDs of stellar atmosphere models predict UV properties of hot Pop III stars that deviate from those of blackbody SEDs scaled to the same effective temperatures and bolometric luminosities . We illustrate this in Fig. 1, where we compare the SEDs of the TLUSTY stellar atmosphere models used in Muspelheim to those of blackbody SEDs for stars at â K. The photon counts per wavelength () differ between blackbody and stellar atmosphere SEDs in all the UV bands, albeit to different extent and in different directions throughout the range of model SEDs shown. It is worth noting, that even though the highest- SED shown ( K, a temperature reached only by rotating Yoon et al. (2012) 20â150 Mâ Pop III stars in our set) exhibits the most blackbody-like shape (with small continuum breaks only), the peak of the stellar-atmosphere SED occurs at shorter wavelengths compared to the corresponding blackbody, an effect previously seen in the stellar atmosphere SEDs for Pop III stars by Bromm et al. (2001) and also for pure hydrogen atmospheres for the very hot central stars of planetary nebulae (Rauch, 2003).
@@
The effect on the ionizing- and LW photon production rates can be seen in Fig. 2 where we compare results generated from blackbody SEDs (dashed lines; in good agreement with the results presented by Murphy et al. 2021b) and stellar atmospheres (solid lines) for the non-rotating Murphy et al. (2021a) stellar evolutionary tracks with initial masses in the range 9â120 Mâ. The general behaviour follows what we would expect from Fig. 1, as the shape111But not the scaling, which in Fig. 1 has been chosen to avoid overlap of the SEDs in the plot of the K, and SEDs in Fig. 1 are representative of early evolution along the , and Murphy et al. (2021a) tracks in Fig. 2. The blackbody SEDs have consistently higher H-ionizing rates, but lower He-ionizing rates, for all the stellar masses plotted, although by no more than -0.3 dex. The behaviour of the He+-ionizing rates is more complex, as the blackbody under-predicts the flux at 85-120 Mâ by -0.3 dex, but then severely starts to over-predict the He+-ionizing flux by up to dex at lower masses due to the presence of a strong continuum break in the stellar atmosphere SEDs at the He+ edge (seen in the K and K SEDs in Fig. 1). In the case of the LW flux, the blackbody SED over-predicts the flux by up to dex at the highest masses (as seen in the K SED), becomes nearly identical at 15 Mâ (as seen in the K SED) and drops below the stellar atmosphere SED by dex by 9 Mâ (as seen in the K SED).
@@
When modelling the impact of Pop III stars on the intergalactic medium, the detailed temporal evolution of the UV photon rates is often neglected in favour of lifetime-integrated quantities (for a counterexample, see Gessey-Jones et al., 2022). This may be reasonable in cases where the lifetimes of the Pop III stars are short compared to timescales related to the thermal evolution of the IGM or changes in cosmic star formation rate density, as is the case for the more massive Pop III stars. In Fig. 3 we present the lifetime-integrated photon production rates of our Pop III models per solar mass, compared to blackbody models from Klessen & Glover (2023) and models for rotating stars with homogeneous chemical evolution from Liu et al. (2025a, further discussed in section 5). For stellar masses , all models except the Yoon et al. (2012) and Liu et al. (2025a) rotating models agree on the H-ionizing and He-ionizing rates to within dex. The Yoon et al. (2012) models at 20 MZAMS 200 Mâ, and all of the Liu et al. (2025a) models, undergo chemically homogeneous evolution which significantly boosts the ionizing photon production of these stars. At lower masses, models based on blackbodies are seen to deviate from those based on stellar atmosphere SEDs by up to dex for H and dex for He.
@@
In the case of the He+-ionizing photon production rates, models based on blackbody SEDs at consistently produce higher values than corresponding models based on stellar atmosphere SEDs, as expected from Fig. 2, with an offset that reaches dex at the lowest masses. Models for rotating Yoon et al. (2012) stars at 20â300 Mâ produce higher He+-ionizing rates than all other models at comparable masses. This increase of He+-ionzing photons due to rotation was already highlighted by Yoon et al. (2012), although the increase compared to non-rotating models becomes much higher in models based on stellar atmosphere SEDs (peaking at dex at 20 Mâ) compared to the blackbody SEDs ( dex) used by Yoon et al. (2012), as shown in their fig. 7. Liu et al. (2025a) is not included in the He+ panel as their spectra are cut of at 124Ă which leaves out a significant portion of the He+-ionizing photons, especially at high Teff (Fig. 1).
@@
For the LW band, models that undergo CHE produce somewhat lower rates (â0.3 dex) than other models at 50 Mâ while lower masses produce rates on the higher side compared with other models. The Klessen & Glover (2023) blackbody models also produce systematically lower rates at low stellar masses (reaching offsets of dex at 1.7 Mâ) compared to models based on stellar atmosphere SEDs. At high masses, the most notable outliers are the models by Schaerer (2002), which at produce much higher rates than all other models (up to dex higher). This discrepancy is discussed further in Section 6.2.
@@ 4 Impact on the the HeII1640 emission line from Population III star clusters
The very high He+-ionizing photon production rates of Pop III rotating Yoon et al. (2012) stars are interesting in the context of HeII emission lines from high-redshift galaxies. In particular, the HeII1640 emission line has been proposed as a potential Pop III signature, as it can become very strong (with rest-frame equivalent widths up to Ă ) in the presence of K stars (e.g. Tumlinson & Shull, 2000; Schaerer, 2002; Raiter et al., 2010; Nakajima & Maiolino, 2022). For non-rotating stars, effective temperatures required to produce Ă are not reached until , which means that a very strong HeII1640 line from an integrated stellar population (Pop III star cluster or Pop III-dominated galaxy) could serve as an indicator of a top-heavy Pop III stellar initial mass function. Recently, a system with Ă and no detectable metal lines, located in the halo of the galaxy GNz-11 at , was found by Maiolino et al. (2024) who interpreted this as evidence of a Pop III IMF extending at least up to 500 Mâ.
@@
In Fig. 4, we explore the temporal evolution of the He+-ionizing rates of our Yoon et al. (2012) models, with and without rotation. Since rotation in these models shifts the early main sequence to lower and slightly lower bolometric luminosities, rotation leads to lower He+ rates during early evolution (although this phase is very brief in the case of the 500 Mâ model). However, since the rotating stars eventually evolve to become hotter and more luminous than their non-rotating counterparts for , the He+-ionizing photon rates of these rotating stars grow by several orders of magnitude at high ages. This effect, coupled to the extended lifetimes of these stars (due to rotational mixing renewing the fuel in the core), explains the boosted time-integrated rates for rotating Yoon et al. (2012) models shown in Fig. 3.
@@
Since Fig. 4 indicates that the He+-ionizing rates of these 20â30 Mâ rotating models can become comparable, for about Myr at the end of their lifetimes, to the peak He+-ionizing rates of 100-150 Mâ stars (maintained over a similar timescale, but during early evolution), it seems that very strong HeII1640 emission in principle could be produced by rotating Pop III stars without the presence of stars.
@@
To explore this, we use the latest version of the photoionization code cloudy (v23.01), last described by Chatzikos et al. (2023) and Gunasekera et al. (2023) to investigate the effect of gas parameters on the line strength and equivalent width of the HeII1640 emission line in gas clouds surrounding single stars from the Yoon et al. (2012) set. We assume primordial gas abundances, a plane-parallel nebula, and an ionization parameter in the range to .
@@
In Fig. 5 we show the resulting equivalent widths of the HeII1640 emission line for a subset of Yoon et al. (2012) tracks. Comparing our results with HeII1640 EWs from Schaerer (2003) and Raiter et al. (2010) â which both assume fully sampled Pop III IMFs â we are able to obtain only slightly higher HeII1640 equivalent widths than the most extreme values seen in their models. However, the major difference is that both Schaerer (2003) and Raiter et al. (2010) require IMFs dominated by very massive stars to reach the high EWs that we are able to obtain for 20 Mâ stars. We also explore other ionization parameters, not shown in the figure, and find that the equivalent widths are largely stable at log(U) -3, but note a significant drop at log(U) = -4 in agreement with Raiter et al. (2010).
@@
Our results indicate that a stellar population of Yoon et al. (2012) rotating stars with an IMF peaking at in principle could produce equally strong HeII1640 equivalent widths as what has been seen for Pop III models with non-rotating stars and more top-heavy IMFs (like the 50-500, Salpeter-slope, and the , , 1â500 Mâ lognormal model from Raiter et al. 2010). However, it is important to note that since the K phase of the 20 Mâ rotating Yoon et al. (2012) model is reached only at the very end of its Myr lifetime, a Pop III system that also hosts more massive stars (with lifetimes as short as Myr for the models considered here) may already have exploded as supernovae and cleared away any remaining gas from the vicinity of the stars, which would lead to very low ionization parameters and reduced HeII1640 equivalent widths. In the case of low-mass haloes and sufficiently powerful supernovae, the gas may even be completely removed from the halo, which would preclude the detection of any emission lines associated with the late evolution of these 20 Mâ stars. Even without supernovae, the initially dense gas around a cluster of Pop III stars in haloes may be expelled to large radii and low densities in response to the radiation field from the stars, with long recombination time-scales and significant leakage of ionizing photons into the intergalactic medium as a result (e.g. Sibony et al., 2022). Detailed simulations would be required to assess whether rotating stars would realistically be able to produce strong HeII emission lines (at the level of the recently detected Maiolino et al. 2024 Pop III galaxy candidate) late in the lifetimes of these stars, even if evolution were to follow the Yoon et al. (2012) models.

### v2

# Ultraviolet photon production rates of the first stars: Impact on the He II 1640 Å emission line from primordial star clusters and the 21-cm signal from cosmic dawn

@@ Abstract
The first stars, the chemically pristine Population III, likely played an important role in heating the intergalactic medium during the epoch of cosmic dawn. The very high effective temperatures ( K) predicted for the most massive Population III stars could also give rise to tell-tale signatures in the emission-line spectra of early star clusters or small galaxies dominated by such stars. Important quantities in modelling their observational signatures include their photon production rates at ultraviolet energies at which photons are able to ionize hydrogen and helium, dissociate molecular hydrogen and cause Lyman- heating. Here, we model the spectral energy distributions of Population III stars to explore how these key quantities are affected by the initial mass and rotation of Population III stars given a wide range of models for the evolution of these stars. Our results indicate that rotating Population III stars that evolve to effective temperatures K could potentially give rise to a very strong HeII 1640 Å emission line in the spectra from primordial star clusters, without requiring stellar masses of indicated by previous models for non-rotating Population III stars. The observable impact on 21-cm signatures from cosmic dawn and the epoch of reionization from our set of rotating stars that evolve to K is modest, except in case of high Population III star formation efficiencies which imprint potentially detectable features in the global 21-cm signal and 21-cm power spectrum.
@@
Stars: Population III – stars: rotation – cosmology: dark ages, reionization, first stars – galaxies: high-redshift1 Introduction
@@
The first generation of stars, the chemically pristine Population III (hereafter Pop III), is believed to have played an important role in heating the early intergalactic medium during the epoch of cosmic dawn, in producing the first elements beyond lithium, and in the production of early black holes (for a recent review, see Klessen and Glover, 2023). Such stars probably started forming around redshift , about 100 Myr after the Big Bang, but may have continued to form in pockets of primordial gas throughout the reionization epoch and well into the post-reionization era (e.g. Johnson, 2010; Mebane et al., 2018; Liu and Bromm, 2020). In terms of observational probes, 21-cm cosmology may be able to probe the nature of Pop III stars at –30 through their impact on the temperature state of the intergalactic medium (e.g. Pochinda et al., 2024; Gessey-Jones et al., 2025; Ventura et al., 2025). At somewhat lower redshifts (–15) one may hope to detect Pop III signatures in individual objects, either among samples of gravitationally lensed, high-redshift stars (Windhorst et al., 2018; Zackrisson et al., 2024), in gravitationally lensed Pop III star clusters and galaxies (Zackrisson et al., 2012; Vanzella et al., 2023; Fujimoto et al., 2025), or within galaxies where Pop III star formation is able to continue alongside metal-enriched star formation (e.g. Sarmento et al., 2018; Venditti et al., 2023; Maiolino et al., 2024; Wang et al., 2024). Other ways of empirically studying the properties of Pop III stars include the chemical signatures of Pop III supernovae in gas absorption systems (e.g. D’Odorico et al., 2023; Sodini et al., 2024; Vanni et al., 2024), in high-redshift quasars (Yoshii et al., 2022) and in second-generation stars in the local Universe (e.g. Fraser et al., 2017; Ishigaki et al., 2018; Jiang et al., 2024; Bonifacio et al., 2025), as well as the direct detection of supernovae (e.g. Moriya et al., 2022; Venditti et al., 2024), gamma-ray bursts (e.g. Lazar and Bromm, 2022), and tidal disruption events (Kar Chowdhury et al., 2024) produced by Pop III stars.
@@
Due to their chemically pristine composition, Pop III stars are expected to be more massive (e.g. Klessen and Glover, 2023), and reach higher main-sequence temperatures (e.g. Schaerer, 2002) than their metal-enriched descendants. This gives rise to higher hydrogen- and helium-ionizing photon production rates compared to metal-enriched stars, and also differences in the Lyman–Werner (LW; -dissociating) and the Lyman-band rates important for the Wouthuysen–Field coupling during cosmic dawn. However, the production of photons in these different energy regimes depends not only on initial mass but also on stellar rotation (e.g. Yoon et al., 2012; Murphy et al., 2021b; Sibony et al., 2022; Liu et al., 2025a, b; Hawcroft et al., 2025), and simulations have indicated that at least some Pop III may rotate at close to break-up speed (Hirano and Bromm, 2018).
@@
Here, we use the Muspelheim (Zackrisson et al., 2024) models for the spectral energy distributions (SEDs) of Pop III stars to explore the impact of different sets of stellar evolutionary tracks on the ultraviolet (UV) radiation output of Pop III stars. These models use stellar atmosphere SEDs instead of the blackbody SEDs adopted in many previous studies (e.g. Yoon et al., 2012; Murphy et al., 2021b; Klessen and Glover, 2023). In section 2, we present these models and in section 3 we report the resulting hydrogen- and helium-ionizing photon production rates, the LW and the Ly-band rates, along with a comparison to results derived from blackbody SEDs. The impact on studies of the UV SEDs of Population III star clusters and galaxies is explored in Section 4 and the impact on 21 cm-cosmology is shown in Section 5. Various future avenues for improvement of the models are discussed in section 6. Section 7 summarises our findings.
@@ 2 SED models of Population III stars
The SED models for Pop III stars (Muspelheim v.1) presented by Zackrisson et al. (2024) couple sets of stellar evolutionary tracks to a grid of stellar atmospheres to produce SEDs as a function of Zero Age Main Sequence (ZAMS) mass and age. Important simplifications used in the Muspelheim v.1 models include the use of stellar atmospheres for spherical stars without mass loss, and with fixed primordial abundances of hydrogen and helium. Hence, these models neglect the non-sphericity that come from high rotation rates, the effects of winds, and the effects of surface helium enhancement and surface self-pollution of metals during the course of evolution. These shortcomings are further discussed in Section 6. In this paper, we make use of Muspelheim SED models based on selected Pop III tracks from Yoon et al. (2012) at –500 M⊙, Windhorst et al. (2018) at –100 M⊙, Murphy et al. (2021a) at –120 M⊙, Volpato et al. (2023) at –1000 M⊙ and Costa et al. (2025) at –200 M⊙. Out of these sets, the Yoon et al. (2012) and Murphy et al. (2021a) sets feature models both with and without rotation. Although Yoon et al. (2012) present stellar evolutionary tracks for several different initial velocities in the range 0.0–0.6 (where is the Keplerian velocity at the equatorial surface), we here limit the discussion to the and 0.4 models, for simplicity.
@@
In this paper, we explore the UV properties of Pop III stars in terms of their production of H-, He- and He+-ionizing photons, as well as their production of LW and Lyman-band photons.
The instantaneous photon production rates , expressed in units of photons per second, are defined as:
where is the monochromatic luminosity (erg-1 s-1 Å-1) at wavelength , is Planck’s constant and the speed of light. The integration limits for the different rates are: and Å for ; and Å for ; and Å for ; Å and Å for ; Å and Å for . We also present , the Lyman-band photon production count per baryon, integrated over the stellar lifetime , given by:
@@
3 Ionizing and Lyman–Werner fluxes
@@
As previously shown by e.g. Bromm et al. (2001), Larkin et al. (2023), and Liu et al. (2025a), the SEDs of stellar atmosphere models predict UV properties of hot Pop III stars that deviate from those of blackbody SEDs scaled to the same effective temperatures and bolometric luminosities . We illustrate this in Fig. 1, where we compare the SEDs of the TLUSTY stellar atmosphere models used in Muspelheim to those of blackbody SEDs for stars at – K. The photon counts per wavelength () differ between blackbody and stellar atmosphere SEDs in all the UV bands, albeit to different extent and in different directions throughout the range of model SEDs shown. It is worth noting, that even though the highest- SED shown ( K, a temperature reached only by rotating Yoon et al. (2012) 20–150 M⊙ Pop III stars in our set) exhibits the most blackbody-like shape (with small continuum breaks only), the peak of the stellar-atmosphere SED occurs at shorter wavelengths compared to the corresponding blackbody, an effect previously seen in the stellar atmosphere SEDs for Pop III stars by Bromm et al. (2001) and also for pure hydrogen atmospheres for the very hot central stars of planetary nebulae (Rauch, 2003).
@@
The effect on the ionizing- and LW photon production rates can be seen in Fig. 2 where we compare results generated from blackbody SEDs (dashed lines; in good agreement with the results presented by Murphy et al. 2021b) and stellar atmospheres (solid lines) for the non-rotating Murphy et al. (2021a) stellar evolutionary tracks with initial masses in the range 9–120 M⊙. The general behaviour follows what we would expect from Fig. 1, as the shape111But not the scaling, which in Fig. 1 has been chosen to avoid overlap of the SEDs in the plot of the K, and SEDs in Fig. 1 are representative of early evolution along the , and Murphy et al. (2021a) tracks in Fig. 2. The blackbody SEDs have consistently higher H-ionizing rates, but lower He-ionizing rates, for all the stellar masses plotted, although by no more than -0.3 dex. The behaviour of the He+-ionizing rates is more complex, as the blackbody under-predicts the flux at 85-120 M⊙ by -0.3 dex, but then severely starts to over-predict the He+-ionizing flux by up to dex at lower masses due to the presence of a strong continuum break in the stellar atmosphere SEDs at the He+ edge (seen in the K and K SEDs in Fig. 1). In the case of the LW flux, the blackbody SED over-predicts the flux by up to dex at the highest masses (as seen in the K SED), becomes nearly identical at 15 M⊙ (as seen in the K SED) and drops below the stellar atmosphere SED by dex by 9 M⊙ (as seen in the K SED).
@@
When modelling the impact of Pop III stars on the intergalactic medium, the detailed temporal evolution of the UV photon rates is often neglected in favour of lifetime-integrated quantities (for a counterexample, see Gessey-Jones et al., 2022). This may be reasonable in cases where the lifetimes of the Pop III stars are short compared to timescales related to the thermal evolution of the IGM or changes in cosmic star formation rate density, as is the case for the more massive Pop III stars. In Fig. 3 we present the lifetime-integrated photon production rates of our Pop III models per solar mass, compared to blackbody models from Klessen and Glover (2023) and models for rotating stars with homogeneous chemical evolution from Liu et al. (2025a, further discussed in section 5). For stellar masses , all models except the Yoon et al. (2012) and Liu et al. (2025a) rotating models agree on the H-ionizing and He-ionizing rates to within dex. The Yoon et al. (2012) models at 20 MZAMS 200 M⊙, and all of the Liu et al. (2025a) models, undergo chemically homogeneous evolution which significantly boosts the ionizing photon production of these stars. At lower masses, models based on blackbodies are seen to deviate from those based on stellar atmosphere SEDs by up to dex for H and dex for He.
@@
In the case of the He+-ionizing photon production rates, models based on blackbody SEDs at consistently produce higher values than corresponding models based on stellar atmosphere SEDs, as expected from Fig. 2, with an offset that reaches dex at the lowest masses. Models for rotating Yoon et al. (2012) stars at 20–300 M⊙ produce higher He+-ionizing rates than all other models at comparable masses. This increase of He+-ionzing photons due to rotation was already highlighted by Yoon et al. (2012), although the increase compared to non-rotating models becomes much higher in models based on stellar atmosphere SEDs (peaking at dex at 20 M⊙) compared to the blackbody SEDs ( dex) used by Yoon et al. (2012), as shown in their fig. 7. Liu et al. (2025a) is not included in the He+ panel as their spectra are cut of at 124Å which leaves out a significant portion of the He+-ionizing photons, especially at high Teff (Fig. 1).
@@
For the LW band, models that undergo CHE produce somewhat lower rates (–0.3 dex) than other models at 50 M⊙ while lower masses produce rates on the higher side compared with other models. The Klessen and Glover (2023) blackbody models also produce systematically lower rates at low stellar masses (reaching offsets of dex at 1.7 M⊙) compared to models based on stellar atmosphere SEDs. At high masses, the most notable outliers are the models by Schaerer (2002), which at produce much higher rates than all other models (up to dex higher). This discrepancy is discussed further in Section 6.2.
@@ 4 Impact on the the HeII1640 emission line from Population III star clusters
The very high He+-ionizing photon production rates of Pop III rotating Yoon et al. (2012) stars are interesting in the context of HeII emission lines from high-redshift galaxies. In particular, the HeII1640 emission line has been proposed as a potential Pop III signature, as it can become very strong (with rest-frame equivalent widths up to Å) in the presence of K stars (e.g. Tumlinson and Shull, 2000; Schaerer, 2002; Raiter et al., 2010; Nakajima and Maiolino, 2022). For non-rotating stars, effective temperatures required to produce Å are not reached until , which means that a very strong HeII1640 line from an integrated stellar population (Pop III star cluster or Pop III-dominated galaxy) could serve as an indicator of a top-heavy Pop III stellar initial mass function (IMF). Recently, a system with Å and no detectable metal lines, located in the halo of the galaxy GNz-11 at , was found by Maiolino et al. (2024) who interpreted this as evidence of a Pop III IMF extending at least up to 500 M⊙ and a total mass of in Pop III stars.
@@
In Fig. 4, we explore the temporal evolution of the He+-ionizing rates of our Yoon et al. (2012) models, with and without rotation. Since rotation in these models shifts the early main sequence to lower and slightly lower bolometric luminosities, rotation leads to lower He+ rates during early evolution (although this phase is very brief in the case of the 500 M⊙ model). However, since the rotating stars eventually evolve to become hotter and more luminous than their non-rotating counterparts for , the He+-ionizing photon rates of these rotating stars grow by several orders of magnitude at high ages. This effect, coupled to the extended lifetimes of these stars (due to rotational mixing renewing the fuel in the core), explains the boosted time-integrated rates for rotating Yoon et al. (2012) models shown in Fig. 3.
@@
Since Fig. 4 indicates that the He+-ionizing rates of these 20–30 M⊙ rotating models can become comparable, for about Myr at the end of their lifetimes, to the peak He+-ionizing rates of 100-150 M⊙ stars (maintained over a similar timescale, but during early evolution), it seems that very strong HeII1640 emission in principle could be produced by rotating Pop III stars without the presence of stars.
@@
To explore this, we use the latest version of the photoionization code cloudy (v23.01), last described by Chatzikos and others (2023) and Gunasekera and others (2023), to investigate the effect of gas parameters on the line strength and equivalent width of the HeII1640 emission line in gas clouds surrounding single stars from the Yoon et al. (2012) set. We assume primordial gas abundances, a plane-parallel nebula, and an ionization parameter in the range to .
@@
In Fig. 5 we show the resulting equivalent widths of the HeII1640 emission line for a subset of Yoon et al. (2012) tracks. Comparing our results with HeII1640 EWs from Schaerer (2003) and Raiter et al. (2010) – which both assume fully sampled Pop III IMFs – we are able to obtain only slightly higher HeII1640 equivalent widths than the most extreme values seen in their models. However, the major difference is that both Schaerer (2003) and Raiter et al. (2010) require IMFs dominated by very massive stars to reach the high EWs that we are able to obtain for 20 M⊙ stars. We also explore other ionization parameters, not shown in the figure, and find that the equivalent widths are largely stable at log(U) -3, but note a significant drop at log(U) = -4 in agreement with Raiter et al. (2010).
Our results indicate that a stellar population of Yoon et al. (2012) rotating stars with an IMF peaking at in principle could produce equally strong HeII1640 equivalent widths as what has been seen for Pop III models with non-rotating stars and more top-heavy IMFs (like the 50-500 , Salpeter-slope model, and the , , 1–500 M⊙ lognormal model from Raiter et al. 2010). Invoking rotating Yoon et al. (2012) Pop III stars to explain the HeII1640 emission in the halo of GNz-11 would not only remove the requirement of Pop III stars to explain the high equivalent width, but also slightly reduce the total mass in Pop III stars required to explain the line luminosity. Because our rotating stars reach somewhat higher ratio than the models used by Maiolino et al. (2024), we find that the total mass in Pop III stars could be reduced by a factor of compared to their estimate, although this would happen only in a scenario where all Pop III stars within the system are rapidly rotating stars caught right in their hottest phase. An IMF strongly peaked around is not favoured by current simulations (for a review, see Klessen and Glover, 2023), but the chemical properties of several extremely metal-poor stars in the Milky Way halo and in ultrafaint dwarfs can be explained by supernovae in the mass range, either from faint supernovae with fallback, hypernovae or a combination of supernovae and winds from rapidly rotating Pop III stars (e.g. Ishigaki et al., 2018; Skúladóttir et al., 2021; Placco et al., 2021; Jeena et al., 2023). Under the assumption of second-generation stars having been enriched by single Pop III supernovae, the exercise of fitting models of supernova ejecta to stellar abundance patterns can give the appearance of a log-normal Pop III IMF peaked around Ishigaki et al. (2018), but folding in duo-enrichment and the propensity for failed supernovae can also make observed abundance patterns consistent with an IMF extending to much lower and higher masses (Jiang et al., 2024). Finally, a Pop III IMF peaked around 20 is also consistent with the Pop III IMF range found by cosmological models of Milky Way formation, which self-consistently follow the chemical evolution across cosmic time and constrain the Pop III IMF by matching the number and properties of observed ancient metal-poor stars (e.g. Koutsouridou et al., 2024).
@@
It is important to note that since the K phase of the 20 M⊙ rotating Yoon et al. (2012) model is reached only at the very end of its Myr lifetime, a Pop III system that also hosts very massive Pop III stars with -–260 (and lifetimes as short as Myr for the models considered here) will already have experienced the explosion of pair-instability supernovae (e.g. Heger and Woosley, 2002). If the hosting halo has low mass, these energetic explosions might have cleared away much of the remaining gas from the vicinity of the stars, which would lead to very low ionization parameters and thus reduced HeII1640 equivalent widths. For the lowest-mass Pop III haloes, a large fraction of the gas may even be removed from the halo (e.g. Bromm et al., 2003; Kitayama and Yoshida, 2005; Mead et al., 2025). This would preclude the detection of any emission lines associated with the late evolution of these 20 M⊙ stars, unless the gas can quickly recollapse (in the Magg et al. 2022b simulations of Pop III haloes, recollapse can happen in as little as Myr, although the median recollapse time is 50 Myr). Even without supernovae, the initially dense gas around a cluster of Pop III stars in haloes may be expelled to large radii and low densities in response to the radiation field from the stars, with long recombination time-scales and significant leakage of ionizing photons into the intergalactic medium as a result (e.g. Sibony et al., 2022). Detailed simulations would be required to assess whether rotating stars would realistically be able to produce strong HeII emission lines (at the level of the recently detected Maiolino et al. 2024 Pop III galaxy candidate) late in the lifetimes of these stars, even if evolution were to follow the Yoon et al. (2012) models.